OpenStudy (ttop0816):

Please help! The range of the function y = secx - 2 is all reals except -3

OpenStudy (anonymous):

What does sec equal to?

OpenStudy (ttop0816):

1/cos!

OpenStudy (anonymous):

What is the range for cosine?

OpenStudy (ttop0816):

it is (-1,1)

OpenStudy (anonymous):

So if that is true what is the range for sec or (1/cos)?

OpenStudy (ttop0816):

the range for sec is (-infinity , -1] U [1 , +infinity)!

OpenStudy (anonymous):

Right so sec cannot be from -1 to 1 correct. Plug those to numbers into your equation to be the numbers that x cannot be.

OpenStudy (ttop0816):

1<y<3??

OpenStudy (anonymous):

y= -1 -2 y=-3 y= 1 -2 y= -1

OpenStudy (ttop0816):

that would be -3<y<-1 correct?

OpenStudy (anonymous):

Yes

OpenStudy (ttop0816):

the overall answer would be -3<y<-1??

OpenStudy (anonymous):

:-} yes

OpenStudy (ttop0816):

could you also check if im correct?? The graph of y = tanx + 1 crosses the y-axis at _____. -1 0 1 <<< (: 2

OpenStudy (anonymous):

Looks good to me